OpenStudy (anonymous):

The area of circle D is 113.1 square meters. The area of sector ADB is 34.6 square meters. Find the Radius of D, The Circumference of D, the measure of arc AB, and the length of the circle the arc does not contain.

OpenStudy (anonymous):

Help if you can

jhonyy9 (jhonyy9):

do you can drawing it please ?

OpenStudy (anonymous):



OpenStudy (anonymous):

looks like that.

jhonyy9 (jhonyy9):

ok so do you know formula of area of a circle ?

OpenStudy (anonymous):

(pi)r^2 Right?

jhonyy9 (jhonyy9):

so make this pi*r^2 = area what you know and what can you calculi from that easy ?

jhonyy9 (jhonyy9):

pi =3,14 - yes ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So work backwards then?

jhonyy9 (jhonyy9):

area what you know is 113,1 unit squares formula area is pir^2 = A 3,14*r^2 =113,1 yes ?

OpenStudy (anonymous):

Yes, so would I divide or subtract 3.14 from 113.1? Seems a bit confusing in this case. After that, I would root the answer?

OpenStudy (anonymous):

For radius, anyway.

jhonyy9 (jhonyy9):

not subtract just divide by 3,14 both sides

jhonyy9 (jhonyy9):

yes for radius - make it and see what you get for radius - yes sure

OpenStudy (anonymous):

After dividing, I got 36.02 and after rooting I got 6.002. Would the circumference then be 36.69?

OpenStudy (anonymous):

Thank you for your help, I just figured it out :D